# Protecting Human Cognition in the Age of AI

### v1

@@ Abstract.
The rapid adoption of Generative AI (GenAI) is significantly reshaping human cognition, influencing how we engage with information, think, reason, and learn. This paper synthesizes existing literature on GenAI’s effects on different aspects of human cognition. Drawing on Krathwohl’s revised Bloom’s Taxonomy and Dewey’s conceptualization of reflective thought, we examine the mechanisms through which GenAI is affecting the development of different cognitive abilities. Accordingly, we provide implications for rethinking and designing educational experiences that foster critical thinking and deeper cognitive engagement and discuss future directions to explore the long-term cognitive effects of GenAI.
conference: 2025; 2025; 2025copyright: none
@@ 1. Introduction
This work explores the impact of GenAI on fundamental aspects of human cognition and how other societal factors, such as heightened stress levels (melchior_work_2007, ) and information overload (woods2002can, ), interact with these effects. While GenAI can both augment and hinder cognitive development, depending on how it is used, this work primarily examines the latter and explores strategies to protect essential cognitive abilities. We focus on examining GenAI’s cognitive effects on novices, such as students, rather than on experts or professionals. However, we also draw some comparisons between the two groups to better understand GenAI’s multifaceted impacts.
We begin by synthesizing the existing literature on the impacts of GenAI on human cognition to delineate the various dimensions of cognition that are being influenced. Building on this, we draw upon Krathwohl’s revised Bloom’s Taxonomy (krathwohl2002revision, ) and Dewey’s seminal work, How We Think (dewey2022we, ) to explore the underlying reasons for the observed effects on human cognition. Based on these insights, we discuss the implications for supporting thinking and learning in the AI era and provide directions for future research to deepen our understanding of the evolving nature of thinking. With this goal, we explore the following research questions:
  • RQ1. What impacts of GenAI on human cognition have been observed, based on current research?
  • RQ2. How is GenAI impacting thinking and learning?
  • RQ3. Accordingly, what are the implications for designing tools to support thinking and learning in the age of AI?
@@ 2. Emerging Cognitive Challenges in the AI Era
GenAI search systems, such as Perplexity.ai, SearchGPT, and Microsoft Copilot, are transforming information seeking behaviors given their ability to synthesize well-structured responses spanning multiple sources. Traditional search engines require active seeking and verification of sources (shah2024envisioning, ), while GenAI can shift users toward passive consumption of potentially one-sided views, diminishing their ability to discern reliable information and limiting exposure to diverse perspectives (venkit2024search, ). The tendency to engage with AI-generated outputs that align with pre-existing beliefs further exacerbates the echo chamber effect (cinelli2021echo, ). Sharma et al. (sharma_generative_2024, ) show that people interacting with a GenAI search system placed selective attention or retention on GenAI responses and spent more time on consonant versus dissonant opinions. Further, the populistic approach of GenAI search systems– prioritizing widely accepted perspectives while sidelining alternative views– can contribute to homogenization, marginalize less visible but valuable information and suppress marginalized identities, languages, cultural practices, and epistemologies (solaiman2023evaluating, ; venkit2024search, ).
@@ 3. Mechanisms of AI-Induced Cognitive Change
To understand how GenAI affects human cognition, we draw upon Bloom’s Revised Taxonomy and Dewey’s theory of reflective thinking. These foundational frameworks provide complementary perspectives on thinking and learning processes and help illuminate the mechanisms through which AI influences cognitive development. Further, we discuss how other societal factors such as widespread stress and anxiety interact with the effects of GenAI on human cognition.
@@ 4. Supporting Thinking & Learning in the AI Era
At the beginning of learning a new skill, cognitive effort and persistence are essential— also conceptualized as ‘productive struggle’ (stanford2025productive, ) or ‘productive failure’ (kapur_productive_2008, ). Similarly, Brown et al. (brown2024cognitive, ) posit building ‘cognitive endurance’ among students, i.e., the ability to sustain effortful mental activity over a continuous stretch of time. This effort, even when it is challenging or frustrating, is needed for deep understanding. Therefore, in the early stages of learning, AI use should be minimal, primarily serving functions such as providing formative feedback (irons2021enhancing, ). This can be implemented by embedding guardrails in educational AI tools that facilitate a gradual increase in learner-AI interactions, ensuring that learners have the ability to exercise independent judgment when they seek AI assistance.
With regard to the cognitive skills in Bloom’s revised taxonomy, the evolving nature of thinking and learning—characterized by increased human-AI collaboration—necessitates a greater emphasis on teaching critical and evaluative skills to ensure effective engagement with GenAI tools and informed decision-making. However, current curricula and high-stakes tests prioritize fostering skills at which AI excels, such as formulaic decision-making (cao2023navigating, ). To address this, test designers should consider emphasizing critical and evaluative skills in institutional and standardized tests, as these directly inform teaching practice (pedulla2003perceived, ). Developing learning activities that require students to actively critique GenAI outputs (singh2024bridging, ; singh2024empowering, ; oates2025chatgpt, ) can also be helpful for fostering the development of such skills.
@@
GenAI chatbots prioritize surface-level discussions over active and deep learning, particularly for novices who often lack prompt engineering skills and foundational understanding of the underlying subject (federiakin2024prompt, ). Building on schema theory (derry1996cognitive, ), which describes how knowledge is organized in the brain as interconnected frameworks or ‘schemas’, integrating cognitive schemas (andersonpearson1984, ) into AI-driven educational tools can help learners identify connections between related ideas and build comprehensive understanding (Rusen2024, ). For such schema-based learning, an example tool could have an interactive interface with a navigable knowledge graph along with a conversational AI agent. When encountering new information in a conversation, the tool could dynamically highlight relevant schemas in the knowledge graph to activate prior knowledge. These features move beyond conversational AI, providing structured scaffolding for deeper learning. Another possibility is enabling learners to highlight parts of GenAI responses, which not only fosters active engagement but also provides feedback to the GenAI system, enabling it to generate more relevant responses to support a learner’s thinking. Finally, introducing small amounts of forced yet desirable engagement with AI-generated outputs—conceptualized as ‘friction’ in human-AI interaction (kazemitabaar2024exploring, )–before learners can use the generated outputs can help promote more active engagement with GenAI tools.
@@
Another key consideration based on Dewey’s concept of reflective thought is to encourage suspended judgment in interactions with GenAI tools. One way to do so is using metacognitive prompts (bannert2013scaffolding, ) which nudge learners to pause and reflect at critical moments–for instance, by encouraging them to consider alternative perspectives or assess their comprehension of GenAI outputs. Metacognitive prompts can help learners be more aware of their thinking, as even those who possess metacognitive knowledge and skills may not always know how to use them spontaneously (bannert2013scaffolding, ). Similarly, recent work has explored the concept of AI-generated ‘provocations’, which highlight the risks, biases, limitations, and alternatives to GenAI recommendations, to encourage critical engagement rather than passive acceptance of AI suggestions (drosos2025makes, ; sarkar2024ai, ).
5. Conclusion & Future Work
This paper examines the impacts of GenAI on human cognition and their underlying mechanisms, providing several key insights that have important implications for education, interaction design, and cognitive development in the AI era. Much of the current research on AI’s cognitive impacts relies on short-term studies, making it difficult to draw definitive conclusions about long-term longitudinal effects on thinking and learning. This presents a pressing need for long-term studies examining how sustained AI use affects cognitive development, particularly for younger users. Moreover, the long-term effects of LLM reasoning frameworks, such as Chain-of-Thought (wei2022chain, ), Tree of Thought (yao2024tree, ), and Graph-of-Thought(besta2024graph, ), on human reasoning need to be studied. Unlike human reasoning, which follows causal chains and implicit thought processes, LLM reasoning is explicit and often deviates from human patterns (Bao2024HowLD, ; Yax2023Studying, ). Whether and how this disparity influences human reasoning warrants further exploration. Next, novel evaluation frameworks need to be developed, particularly in educational settings, to assess thinking and learning–extending beyond conventional metrics to incorporate dimensions such as depth of learning, cognitive engagement, critical thinking, and the development of metacognitive skills. Finally, further research is needed to examine how cultural and socioeconomic factors shape the cognitive effects of AI use, with a focus on access disparities, digital literacy, and cultural variations in instructional approaches.

### v2

@@ Abstract.
The rapid adoption of Generative AI (GenAI) is significantly reshaping human cognition, influencing how we engage with information, think, reason, and learn. This paper synthesizes existing literature on GenAI’s effects on different aspects of human cognition. Drawing on Krathwohl’s revised Bloom’s Taxonomy and Dewey’s conceptualization of reflective thought, we examine the mechanisms through which GenAI is affecting the development of different cognitive abilities. We focus on novices, such as students, who may lack both domain knowledge and an understanding of effective human-AI interaction. Accordingly, we provide implications for rethinking and designing educational experiences that foster critical thinking and deeper cognitive engagement.
conference: CHI2025 Workshop on Tools for Thought; April 26, 2025; Yokohama, Japancopyright: none
@@ 1. Introduction
This work explores the impact of GenAI on fundamental aspects of human cognition and how other societal factors, such as heightened stress levels (melchior_work_2007, ) and information overload (woods2002can, ), interact with these effects. While GenAI can both augment and hinder cognitive development depending on how it is used, this work primarily examines the latter and explores strategies to protect essential cognitive abilities. We focus on examining GenAI’s cognitive effects on novices, such as students, rather than on experts or professionals. However, we also draw some comparisons between the two groups to better understand GenAI’s multifaceted impacts.
We begin by synthesizing the existing literature on the impacts of GenAI on human cognition to delineate the various dimensions of cognition that are being influenced. Building on this, we draw upon Krathwohl’s revised Bloom’s Taxonomy (krathwohl2002revision, ) and Dewey’s seminal work, How We Think (dewey2022we, ) to explore the underlying reasons for the observed effects on human cognition. Accordingly, we propose adopting Dewey’s conceptualization of critical thinking—as reflective thinking—to both evaluate the impact of GenAI on learners’ critical thinking and inform the design of interventions aimed at fostering critical thinking. Further, we provide insights on applying Dewey’s framework to support the evolving nature of thinking and learning.
@@ 2. Emerging Cognitive Challenges in the AI Era
GenAI search systems, such as Perplexity.ai, SearchGPT, and Microsoft Copilot, are transforming information seeking behaviors given their ability to synthesize well-structured responses spanning multiple sources. Traditional search engines require active seeking and verification of sources (shah2024envisioning, ), while GenAI may shift users toward passive consumption of potentially one-sided views, diminishing their ability to discern reliable information and limiting exposure to diverse perspectives (venkit2024search, ). The tendency to engage with AI-generated outputs that align with pre-existing beliefs further exacerbates the echo chamber effect (cinelli2021echo, ). Sharma et al. (sharma_generative_2024, ) show that people interacting with a GenAI search system placed selective attention or retention on GenAI responses and spent more time on consonant versus dissonant opinions. Further, the populistic approach of GenAI search systems– prioritizing widely accepted perspectives while sidelining alternative views– can contribute to homogenization, marginalize less visible but valuable information and suppress marginalized identities, languages, cultural practices, and epistemologies (solaiman2023evaluating, ; venkit2024search, ).
@@ 3. Mechanisms of AI-Induced Cognitive Change
To understand how GenAI affects human cognition, we draw upon Bloom’s Revised Taxonomy and Dewey’s conceptualization of reflective thinking. These foundational frameworks provide complementary perspectives on thinking and learning processes and help illuminate the mechanisms through which AI influences cognitive development. Further, we discuss how other societal factors such as widespread stress and anxiety interact with the effects of GenAI on human cognition.
@@ 4. Supporting Thinking & Learning in the AI Era
Using Bloom’s Revised Taxonomy, we explored how the cognitive processes that support learning are being impacted in contexts mediated by GenAI. Our analysis reveals that the evolving nature of thinking and learnin-g—characterized by increased human-AI interactions—necessitates a greater emphasis on teaching critical and evaluative skills to ensure effective engagement with GenAI tools. However, current curricula and high-stakes tests prioritize fostering skills at which AI excels, such as formulaic decision-making (cao2023navigating, ). To address this, test designers should consider emphasizing critical and evaluative skills in institutional and standardized tests, as these directly inform teaching practice (pedulla2003perceived, ). Developing learning activities that require students to actively critique GenAI outputs (singh2024bridging, ; singh2024empowering, ; oates2025chatgpt, ) can also be helpful for fostering the development of such skills.
At the beginning of learning a new skill, cognitive effort and persistence are essential— also conceptualized as ‘productive struggle’ (stanford2025productive, ) or ‘productive failure’ (kapur_productive_2008, ). Similarly, Brown et al. (brown2024cognitive, ) posit building ‘cognitive endurance’ among students, i.e., the ability to sustain effortful mental activity over a continuous stretch of time. This effort, even when it is challenging or frustrating, is needed for deep understanding. Therefore, in the early stages of learning, AI use should be minimal, primarily serving functions such as providing formative feedback (irons2021enhancing, ). This can be implemented using guardrails in educational AI tools that facilitate a gradual increase in learner-AI interactions, ensuring that learners have the ability to exercise independent judgment when they seek AI assistance.
@@
GenAI influences cognitive development in complex and multifaceted ways, with its impact largely dependent on who engages with the technology and how it is used. In particular, it is passive engagement and over-reliance on AI that pose significant risks to users’ thinking. This calls for the development of novel evaluation frameworks to assess learning, particularly in educational contexts—moving beyond traditional metrics to assess dimensions such as cognitive and metacognitive engagement, critical thinking, and depth of learning. We propose using Dewey’s prerequisites for reflective thinking as a framework for both: (i) identifying passive AI use based on unmet prerequisites of reflective thinking in human-AI interactions, and (ii) designing interventions to foster critical thinking. We now illustrate several applications of this framework.
The first prerequisite—a state of perplexity, confusion, or doubt prompting inquiry—requires learners to reflect on gaps in their understanding, a process that may be undermined by the immediacy of AI-generated responses. To support this pre-requisite, one approach is to enable learners to highlight parts of GenAI responses. This could also provide feedback to the GenAI system, enabling it to generate more relevant responses to support a learner’s thinking. Similarly, introducing small amounts of forced yet desirable engagement with AI-generated outputs—conceptualized as ‘friction’ in human-AI interaction (kazemitabaar2024exploring, )—before learners can use the generated outputs can help promote more active engagament.
@@
Next, GenAI chatbots may prioritize surface-level discussions over persistent exploration of ideas, which is a another prerequisite of reflective thinking. This tendency may be especially pronounced for novices, who often lack the skills for effective prompt engineering and a foundational understanding of the subject matter (federiakin2024prompt, ). Building on schema theory (derry1996cognitive, ), which describes how knowledge is organized in the brain as interconnected frameworks or ‘schemas’, integrating cognitive schemas (andersonpearson1984, ) into educational tools can facilitate learners’ ability to identify connections between related ideas and link them to prior knowledge, thereby fostering comprehensive understanding (Rusen2024, ). For such schema-based learning, an example tool could have an interactive interface with a navigable knowledge graph along with a conversational AI agent. When encountering new information, the tool could dynamically highlight relevant schemas in the knowledge graph to activate prior knowledge.
@@
For suspended judgment—a prerequisite for reflective thinking undermined by GenAI’s persuasiveness and alignment with users’ preexisting beliefs—metacognitive prompts offer a promising intervention (lin1999supporting, ). These prompts nudge learners to pause and reflect at critical moments, for instance, by encouraging them to consider alternative perspectives or assess their comprehension of GenAI outputs. Metacognitive prompts can help learners be more aware of their thinking, as even those who possess metacognitive knowledge and skills may not always know how to use them spontaneously (bannert2013scaffolding, ). Similarly, recent work has explored the concept of AI-generated ‘provocations’, which highlight the risks, biases, limitations, and alternatives to GenAI recommendations (drosos2025makes, ; sarkar2024ai, ).
5. Conclusion
This paper examines the impacts of GenAI on human cognition, providing several key insights that have important implications for education, interaction design, and cognitive development in the AI era. Much of the current research on AI’s cognitive impacts relies on short-term studies, making it difficult to draw definitive conclusions about long-term longitudinal effects. This presents a pressing need for long-term studies on how sustained AI use affects cognitive development, particularly for younger users who are most vulnerable to the negative effects of GenAI overuse.